Pittman Company is a small but growing manufacturer of telecommunications equipment. The company has no sales force of its owComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute PittmaComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Assume that PiComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Determine theComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the de

Please show the excel work for the question.

Thank you
Answer #1
Solution 1:
15% commission 20% commission own sales force
Sales 17000000 17000000 17000000
Variable expenses:
Manufacturing 7650000 7650000 7650000
Commission 2550000 3400000 1275000
Total variable expenses 10200000 11050000 8925000
Contribution margin 6800000 5950000 8075000
Variable expenses ratio 60.00% 65.00% 52.50%
Contribution margin ratio 40.000% 35.000% 47.500%
Fixed expenses:
Manufacturing overhead 2380000 2380000 2380000
Marketing 119000 119000 2669000
Administrative 1840000 1840000 1761800
Interest 595000 595000 595000
Total fixed expenses 4934000 4934000 7405800
Income before income taxes 1866000 1016000 669200
Income taxes(30%) 559800 304800 200760
Net income 1306200 711200 468440
Break Even Point in dollars (Total Fixed Expenses/Controbution margin ratio) 12335000 14097143 15591158
Solution 2:
Income before tax as per income statement 1866000
Total Fixed Expenses (when 20% commission) 4934000
Total 6800000
/ Contribution margin ratio (When 20% commission) 35.000%
Sales required to generate target profit 19428571
Solution 3:
Let Sales Volume be "X"
Total Cost when commission 20% = Total Cost for Own sales Force
0.6500*X +4934000 = 0.5250*X +7405800
0.125*X = 2471800
X = 2471800/0.125
X = 19774400
Hence, Sales volume required is $19,774,400
Solution 4:
15% commission 20% commission own sales force
Contribution margin 6800000 5950000 8075000
Income before taxes 1866000 1016000 669200
Degree of Operating Leverage (Contribution margin/Income before taxes) 3.64 5.86 12.07